- Press Release
- Sep 29, 2022
NASA Space Station On-Orbit Status 13 May 2004
All ISS systems continue to function nominally, except those noted previously or below.
Both crewmembers spent about 3.5 hours on the lengthy and demanding task of transferring disposable cargo to 13P/Progress-260 at the Service Module (SM) aft end, supported by IMS (inventory management system) barcode tracking, and loading it onboard according to previously uplinked instructions. [The disposal manifest comprises 84 different equipment types, including filled waste bags, discarded hardware parts, used-up batteries, foam material, etc. Instructions emphasize the need for proper restraining of the stowed items to prevent shifting during 13P departure on 5/24.]
CDR Padalka performed the regular once-a-week maintenance power-cycle (reboot) on the operational PCS laptops and also restarted the U.S. OCA comm router laptop (every two weeks).
Gennady worked at the Lab CHeCS (crew health care systems) rack to perform a newly developed one-hour calibration check on the onboard U.S. RSP (Respiratory Support Pack) to determine the unit’s performance for timing (breaths per minute) and volumetric flow. Results will be used for hardware integrity verification and certification life extension, ensuring that an afflicted crewmember receives sufficient oxygen. [After several cracks were found on the ground in the autovent/control module of two RSPs, the new procedure was developed for testing the on-board RSP for performance, since it cannot be visually inspected for cracks.]
FE/SO Michael Fincke changed the lens on the EarthKAM system’s Kodak ESC 460C electronic still camera at its SM window, going from 50mm to the 180mm-lens configuration. [61 schools & over 3,680 students are currently using EarthKAM (EK). The students are hard at work on Day 2, selecting targets and submitting their image requests. Over 400 images have been requested so far, 245 have been downlinked and posted on the EK website ( http://www.earthkam.ucsd.edu ). EarthKAM deactivation and stow is scheduled for 5/15 (Saturday).]
Mike conducted the daily leak check of the Lab window’s inter-pane space (“Volume D”), using the “Aeolus” scopemeter with pressure probe. Today’s pressure reading was 465 Torr (mmHg), or 9 psi.
Both crewmembers again had 60 minutes for general station familiarization and adaptation, as is standard daily rule for the first two weeks after starting station residency.
Fincke completed the daily routine maintenance of the SM’s SOZh life support system, including the routine replacements in the ASU toilet system, and Padalka prepared the daily “delta” file for updating the IMS (inventory management system) database(s).
The crew completed their daily 2.5-hr. physical exercise program on TVIS, RED exerciser and VELO cycle with load trainer. [Engineers continue to investigate the crew-reported problems with the TVIS control panel (resets to zero in motorized mode). While diagnostics is underway, Mike and Gennady are using the treadmill in passive mode, but with VIS (vibration isolation system) active.]
At about 10:30am EDT, the crew had their regular weekly teleconference with the JSC Astronaut Office (Kent Rominger), via S-band S/G-2.
Also at 10:30am, the ground-commanded third round of the four-day transitioning of four MDM computers on the S0 truss element to new software (s/w) began. [Activities today consisted of powering up the backup EXT MDM (EXT-2) and loading it with s/w version R4, then swapping it with the primary EXT-1 computer, which now is the backup MDM. Tomorrow, it too will be transitioned to R4.]
Update on RPCM R&R EVA: For replacing the failed RPCM (remote power controller module) during the EVA on 6/10, the power assembly SPDA (secondary power distribution assembly) on the port side must be accessed. An issue has come up with structural interference between the SPDA door and the Node 2 fluid umbilical tray, which prevents the door from opening enough to complete the RPCM replacement. The problem was known pre-launch but was not corrected on the ground since the analysis focused only on the “ISS Complete” configuration, i.e., with Node 2 installed. Clearing the interference (by removing launch restraint bolts on the fluid umbilical structure and a stanchion bolt [partially] to allow the tray to be pulled enough [~1 inch] out of the way for the door to slide open) may greatly change EVA content and timeline. Investigation continues, and the uplink of the complete EVA timeline procedures scheduled for yesterday was deferred until necessary steps to resolve the issue are defined.
Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Tropical Depression, Indian Ocean (Dynamic event. Predicted to be under track in the west-central Indian Ocean, this storm was designated Tropical Storm 23 but it weakened. It is now strengthening again and is well formed. It should provide dramatic views in the low light of a setting sun), Johannesburg, South Africa (long view of the entire Witwatersrand, looking left. The Witwatersrand is a string of cities, centered on Johannesburg. Originally based on gold mining, manufacturing has taken over as the main employer and wealth generator, making this region the prime economic center in Africa south of the Sahara), and Patagonian Glaciers (two passes; weather is improving for views of both the dry [east, Patagonian] side of the Andes as well as the usually cloudy west [Pacific] side. Crew was to try for the small glaciers. An oblique context view will help), (and thanks for your efforts in getting context views thus far).
CEO images can be viewed at the websites.
See also the website “Space Station Challenge” at
U.S. and Russian Segment Status (as of today, 12:28pm EDT).
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
- Elektron O2 generator is On. Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is Off. TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20). SFOG slot #2 fan suspect (not usable).
- SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 25.3; ppO2 (mmHg) — 157.4; ppCO2 (mmHg) — 3.2;
- SM Transfer Compartment: Pressure (mmHg) — 751; temperature (deg C) — 20.7.
- FGB Cabin: Pressure (mmHg) — 756; temperature (deg C) — 21.7.
- Node: Pressure (mmHg) — 748.34; temperature (deg C) — 22.7 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- U.S. Lab: Pressure (mmHg) — 750.47; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- Joint Airlock (Equip. Lock): Pressure (mmHg) — 750.57; temperature (deg C) — 25.3; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- (n/a = data not available)
- PMA-1: Shell heater temp (deg C) — n/a
- PMA-2: Shell heater temp (deg C) — n/a.
Electrical Power Systems (EPS):
- Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Directed position (blind, dual-angle mode with “night glider”/”sun slicer” drag reduction sweep, non solar-tracking” 2B @ 235 deg, 4B @ 125 deg.)
- SM batteries: Battery #6 is off line; battery #7 is in Cycle (RTs) mode; all other batteries (6) are on line in “Partial Charge” mode.
- FGB batteries: Battery #2 is off line; battery #4 is in Cycle (RTs) mode; all other batteries (4) are on line in “Partial Charge” mode.
- Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.
Command & Data Handling Systems:
- C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
- GNC-2 MDM is prime; GNC-1 is backup.
- INT-2 is operating; INT-1 is Off.
- EXT-1 is On (primary), EXT-2 is Off.
- LA-1, LA-2 and LA-3 MDMs are all operating.
- PL-1 MDM is Off; PL-2 MDM is Operational.
- APS-1 (automated payload switch #1) and APS-2 are both On.
- SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
- SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
- FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.
- Total propellant load available: 3725 (8212 lb) as of 5/5/04; [SM(693) + FGB(2373) + Progress M-1(659)]. (Capability: SM — 860 kg; FGB — 6120 kg).
Attitude Control Systems:
- 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04).
- State vector source — U.S. SIGI-1 (GPS)
- Attitude source — U.S. SIGI-1 (GPS)
- Angular rate source — RGA-1
- LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.
Communications & Tracking Systems:
- FGB MDM-1 is powered Off; FGB MDM-2 is operational.
- All other Russian communications & tracking systems are nominal.
- S-band is operating nominally (on string 2).
- Ku-band is operating nominally (may require a mask).
- Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
- Video subsystem operating nominally.
- HCOR (high-rate communications outage recorder) is operating nominally.
- SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
- MBS: KA (keep alive) power on both strings.
- MT: latched and mated at WS4.
- POA: KA power on both strings.
- RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.
ISS Orbit (as of this morning, 6:45am EDT [= epoch]):
- Mean altitude — 360.7 km
- Apogee — 368.1 km
- Perigee — 353.2 km
- Period — 91.76 min.
- Inclination (to Equator) — 51.63 deg
- Eccentricity — 0.0011012
- Orbits per 24-hr. day — 15.69
- Mean altitude loss last 24 hours — 75 m
- Revolutions since FGB/Zarya launch (Nov. ’98) — 31286
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html